OpenStudy (anonymous):

PLEASE HELP ME! Final t0day! (1)A poster is to have an area of 180in^2with 1 inch margins at the bottom and sides and a 2 inch margin at the top. What dimension will give the largest printed area?

OpenStudy (anonymous):

call length of poster x and width y. then you knowo \[xy=180\] making \[y=\frac{180}{x}\] area of printed part is \[(x-2)(y-4)=(x-2)(\frac{180}{x}-4)=\frac{4(45-x)(x+2)}{x}\]

OpenStudy (anonymous):

This is an optimizastion problem: Maximize: A = (w-2)(h-3) Constraint: 180=wh

OpenStudy (anonymous):

derivative if \[\frac{360-4x^2}{x^2}\] set equal zero, get \[x=3\sqrt{10}\] check my algebra

OpenStudy (anonymous):

psycho i right i am wrong. i should learn to read

OpenStudy (anonymous):

it says one inch around and two at the top. so equation i had was incorrect.

OpenStudy (anonymous):

should be \[A(x)=(x-2)(\frac{180}{x}-3)=-3x-\frac{360}{x}+186\]

OpenStudy (anonymous):

solve the constrint for h and plug it into the equation for the printed area. A(w)=(w-2)(180/w - 3)

OpenStudy (anonymous):

\[A'(x)=-3+\frac{360}{x^2}\] set equal zero get \[3x^2=360\] \[x=2\sqrt{30}\]

OpenStudy (anonymous):

\[A(w)=186-3w-\frac{360}{w}\] \[ A'(w)=-3+\frac{360}{w^2}\] From the derivative you can find three critical points w=0, \[w=\pm \sqrt{\frac{360}{3}} = \pm10.9544512\] since we are working with dimensions we know w>0, so only 10.95445 makes sense.

OpenStudy (anonymous):

So we know that the maximum printed area will be at w=10.95445

OpenStudy (anonymous):

h= 180/10.95445 = 16.4316767